OpenStudy (anonymous):

Solve: (1/3)(6 – 3t) + 18 = 0

OpenStudy (anonymous):

(1/3)(6-3t)+18 = 0 (0.33)(6 - 3t)+18 = 0 (0.33)(6-3t) = -18 Does this help any?

OpenStudy (anonymous):

no man